OpenStudy (christos):

derivative of f(x)=(2x + 1)^3

OpenStudy (anonymous):

Use the chain rule: \[f'(x) = 3(2x+1)^2*2\] So , take derivative considering the thing inside the parenteses as a n "x" and then multiply by the derivative of the inside thing, in this case 2

mathslover (mathslover):

First of all, I have : power = 3 and x = 2x +1 so : \(3(2x+1)^2\) Now, differentiate (2x+1 ) = 2 + 0 = 2 so I have : \(3(2)(2x+1)^2\) = \(6(2x+1)^2

mathslover (mathslover):

Let it be : \(f(x) = X^3\) where \( X = 2x + 1\) So : differentiation of \(X^3 = 3X^2\) Now differentiate X : \(X = 2x + 1\) So , derivative of 2x + 1 (which is equal to X) , \(2 + 0\) = \(2\) So I will write this as : \(3X^2 * 2 = 6X^2 = 6(2x+1)^2\)

mathslover (mathslover):

@Christos , got it?

mathslover (mathslover):

Formula : Let : \(y = ( f(x) )^n \) \(\boxed{\large{\cfrac{dy}{dx} = n (f(x))^{n-1} * \cfrac{d(f(x))}{dx} }}\) Rule used : Chain rule.

mathslover (mathslover):

@Christos , you there?

mathslover (mathslover):

If you're unable to get this method : I am giving you another question just like yours one. \(\LARGE{\textbf{Example : }}\) \(\large{\textbf{Differentiate : } \quad (3x+1)^2 }\) \(\LARGE{\textbf{Solution : }} \) Now, let , \(\large{ y = (3x+1)^2 }\) As per the formula I have : \(\boxed{\large{\cfrac{dy}{dx} = n (f(x))^{n-1} * \cfrac{d(f(x))}{dx} }}\) So : \(\large{\cfrac{dy}{dx} = 2 ( 3x+1 ) ^{2 - 1} * \cfrac{d(3x+1)}{dx}} \\ \implies \cfrac{dy}{dx } = 2(3x+1) * (3x^{1-1} + 0 ) \\ \implies \cfrac{dy}{dx} = 2(3x+1) * 3 \\ \\ \\ \boxed{\cfrac{dy}{dx} = 6(3x+1) } ---- [\textbf{Answer} ] \)

OpenStudy (christos):

I got it thanks!

mathslover (mathslover):

Great to know, @Christos . Best of luck.